a) The trait for eye color in a population has two alleles at one locus. The frequency of the homozygous dominant is 0.1521. If the population is in Hardy-Weinburg equilibrium, what is the frequency of the recessive allele in the population?

b) The trait for lactose tolerance in a population has two alleles at one locus. The lactose tolerance trait is dominant and the lactose intolerance is recessive. The frequency of the lactose tolerance phenotype is 0.75. If the population is in Hardy-Weinburg equilibrium, what is the frequency of the recessive allele in the population?

c) The trait for flower width in a population has two alleles at one locus. The wide flowers trait is dominant and the narrow flowers is recessive. The frequency of the recessive allele is 0.32. If the population is in Hardy-Weinburg equilibrium, what is the predicted frequency of the wide flowers phenotype in the population?
Answer #1

phenotypic frequency of dominant phenotype= genotypic frequency of heterozygous + genotypic frequency of homozygous dominant

alleleic frequency of recessive allele= square root of genotypic frequency of homozygous recessive

a) sum of allelic frequencies= 1

here there are 2 alleles so

the frequency of dominant allele+ frequency of recessive allele=

frequency of recessive allele =1-frequency of the dominant allele

frequency of dominant allelle= square root of the genotypic frequency of homozygous dominant

= square root of 0.1521

= 0.39

so frequency of recessive allele = 1-0.39

= 0.61

b) sum of phenotypic frequencies= 1

frequency of lactose intolerance + frequency lactose tolerance= 1

frequency of lactose intolerance = 1- frequency of lactose tolerance

= 1- 0.75

= 0.25

so allelic frequency of lactose intolerance= square root of genotypic frequency of lactose intolerance

= square root of 0.25

= 0.5

so frequency of recessive allele= 0.5

c) frequency of the recessive allele is 0.32

frequency of narrow flowers =( frequency of narrow allele)^2

= ( 0.32)^2

= 0.1024

sum of phenotypic frequencies= 1

Frequency of narrow flowers+ frequency of wide flowers=1

frequency of wide flowers = 1- Frequency of narrow flower

= 1- 0.1024

= 0.8976